Sum of three numbers in an AP is 15 and sum of squares of its 1st and 3rd term is 58. Find the numbers

Grade:7

1 Answers

Susmita
5 years ago
Let the numbers be
(a-r),a,(a+r).
So
(a-r)+a+(a+r)=15
Or,3a=15
Or,a=5
Also
(a-r)2+(a+r)2=58
Or,a2-2ar+r2+a2+2ar+r2=58
Or,a2+r2=29
or,r2=29-52=4
or,r=2
So the numbers are
3,5,7
If it helps please approve.